how to find the angle between the pair of straight lines x^2-2xy+y^2?

Grade:10

1 Answers

Umakant biswal
7 years ago
@ mukilan 
i can give u the correct approach for this question . kindly solve that 
the two equation will be given to you in the question , 
step 1 = try to find out the slope of the line by comparing with the equation , y = mx + c 
suppose i am taking an example 
two equation are given as 
2x + 3y -1 = 0 
3y= -2x + 1 
y = -2/3 x + 1/3 
by comparing the equation y = mx + c 
m1 = – 2/3 
3x-2y+8 = 0 
2y = 3x + 8 
y = 3/2 x + 4 
slope m2 = 3/2 
angle between two lines , tan theta = {m2-m1 } / 1+m1m2 = 11/0 
so, theta = pie / 2 
angle between two lines = 90 degree 
HOPE IT CLEARS YOUR DOUBT 
KINDLY APPLY THE SAME METHOD TO THIS QUES ALSO , 
ALL THE BEST ..